# RAT [Uncaptioned image]: Bridging RNN Efficiency and Attention Accuracy in Language Modeling

### v1

@@ B.1 Latency
To supplement Fig. 3 in main text, we put the concrete latency number of a single layer in Table 6, Table 7, and Table 8.
Table 6: Single layer (including input and output projections) training time across different sequence lengths. The latency (ms) is tested on 200K tokens.
Model 4096 8192 16384 32768 65536 131072 262144 RNN 77.00 78.42 77.50 83.50 115.83 130.62 195.30 RAT(L=128) 150.19 150.76 156.65 154.52 159.94 165.07 206.46 RAT(L=64) 146.25 148.46 151.10 155.83 165.58 177.78 227.76 RAT(L=16) 150.08 153.63 158.98 173.89 202.00 249.79 378.11 Attention 122.06 176.44 277.61 474.90 877.14 1713.82 3417.48 Attention/RAT(L=16) 0.81$\times$ 1.15$\times$ 1.75$\times$ 2.73$\times$ 4.34$\times$ 6.86$\times$ 9.04$\times$
Table 7: Single layer (including input and output projections) prefilling time across different sequences lengths. The latency (ms) is tested on 200K tokens.
Model 4096 8192 16384 32768 65536 131072 262144 RNN 24.93 24.75 25.32 27.09 30.54 39.40 56.16 RAT(L=128) 44.74 44.78 45.09 45.50 46.67 48.62 51.75 RAT(L=64) 43.46 44.27 44.27 44.83 47.08 50.83 57.38 RAT(L=16) 44.03 44.90 45.53 51.08 57.31 71.18 99.53 Attention 36.93 52.71 80.21 135.68 245.14 494.62 997.50 Attention/RAT(L=16) 0.84$\times$ 1.17$\times$ 1.76$\times$ 2.66$\times$ 4.28$\times$ 6.95$\times$ 10.02$\times$
Table 8: Single layer (including input and output projections) generation time at the specified position. The latency (ms) is tested on generating batches of tokens with $B=64$, $B=512$, and $B=1024$.
@@ B.2 Maximum throughput
As RAT also reduces cache memory usage, we report the maximum throughput of the full 1.3B model across different sequence lengths in Fig. 5. For example, RAT(L=16) achieves an approximately 10.2×10.2$\times$10.2 × higher throughput than the baseline attention model at $T=4096$, and improves further to 15.6×15.6$\times$15.6 × at $T=16384$. Similarly, RAT(L=16)-SWA reaches 4.5×4.5$\times$4.5 ×, 8.0×8.0$\times$8.0 ×, and 13.9×13.9$\times$13.9 × higher throughput over attention at $T=4096$, $8192$, and $16384$, respectively.
Figure 5: We measure the maximum throughput of the full 1.3B model for generating 1024 tokens under different prefilling lengths. For each total sequence length $T$, the prefilling length is set to $T-1024$. For example, $T=4096$ corresponds to a prefilling of 3072 tokens, while $T=8192$ and $T=16384$ correspond to 7168 and 15360 tokens, respectively. When the sequence length increases, the maximum throughput ratio between RAT and attention rises from 10.2×10.2$\times$10.2 × to 15.6×15.6$\times$15.6 ×, highlighting the strong efficiency advantage of RAT in long-context generation.
@@ Appendix C Accuracy
C.1 Reversed design of RAT
Instead of employing recurrence within each chunk, we also explored a reversed design that applies attention inside the chunk, followed by recurrence to capture long-range dependencies. For this reversed design, standard RoPE is sufficient. However, with respect to parameter allocation, we found it crucial to retain the key vector in the attention module as a full tensor, as sharing it together with the query vector leads to a collapse into single-head attention. Thus, our final reversed design shares the query vector across attention heads, employs low-rank matrices for the projection used for output gate, and preserves full projections for both the forget gate and the key vector. Results of the reversed design are put in Table 9. It can be observed that, under identical FLOP constraints, this design significantly underperforms RAT. For example, to match the FLOPs budget of $\mathcal{O}(T/64)$, it requires chunk sizes of 128, 256, and 512 for sequence lengths of 8192, 16384, and 32768, respectively. Most of the computational cost is spent within a small local window of 128 tokens, while excessive burden is placed on the inter-chunk recurrence, ultimately leading to high perplexity. Only when we reduce the number of chunks to 16 does the perplexity drop below 14.00. Therefore, while the reversed design also allows for interpolation between attention and RNN (reducing to attention when $L=T$, and to RNN when $L=1$), we choose to focus on RAT due to its more efficient utilization of FLOPs.
@@ C.2 Length generalization
Because of the use of softmax-based attention and RoPE at the inter-chunk level, it is reasonable to expect that RAT may also face challenges in length generalization, where the model is trained on short sequences but evaluated on much longer ones. To study this, we consider SWA variants using either RoPE or NoPE for attention and RAT, following recent practices that interleave attention with local attention modules and apply positional encodings only in the local attention. This design has been adopted in practice and has shown promising results [51].
@@
In conclusion, we find that RAT can outperform the attention module in length generalization, both with RoPE and with NoPE. Still, the problem is not fully solved, as none of the models reach the stability level of the simple RNN. While many techniques have been proposed to improve length extrapolation in attention models, they should also be considered for RAT, as RAT essentially redirects attention from every position to inter-chunk locations. For example, RoPE extrapolation in the attention module can be improved using methods from Peng et al. [52], bloc97 [53], and NoPE extrapolation has been studied in Wang et al. [54], which points out that NoPE still has a context length limit, though it performs better than RoPE. Their work links the failure to shifts in attention distributions and proposes tuning the temperature of attention heads to improve extrapolation. We believe such methods could also be applied to RAT, and leave further exploration of these techniques to future work.
C.3 16K context window
Table 11: Performance of 1.3B models with a 16K context window, including pretraining perplexity, commonsense reasoning accuracy (short-context), and SFT results on long-context tasks. For SFT, we focus on tasks that still have long context even with the 16K context window, such as QMSum and NarrativeQA3. For NarrativeQA3, we adopt the passage-only setting from the original paper instead of the summary+passage setup used in NarrativeQA2. Although passages are still truncated at 16K, most information are preserved, making passage-only a more meaningful and challenging evaluation setting.
Model Throughput Pretrain Evaluation SFT
token/sec Val. ARC-C ARC-E HellaSwag LAMBADA PIQA WinoGrande Avg. NarrativeQA3 QMSum
RNN 130K 8.09 37.4 65.7 54.2 37.6 72.6 53.4 53.5 15.4 22.8
RAT(L=16) 10376 7.52 36.5 64.8 55.8 45.6 71.5 53.5 54.6 21.6 24.2
Attention 667 7.44 36.8 67.5 57.0 43.9 73.2 57.1 55.9 23.3 23.8
Attention-SWA 1347 7.45 36.9 65.8 56.6 47.6 72.5 55.6 55.8 20.3 23.3
RAT(L=16)-SWA 9244 7.42 38.5 67.5 57.6 47.7 73.3 57.0 56.9 22.6 24.6
We also pretrain the models with a 16K context length. However, we observe that directly pretraining with such a long context window may lead to issues, as most web data is relatively short and needs to be concatenated by the separate token to fill the window. Due to the lack of naturally long-context data, the model may fail to learn true long-context capabilities. We believe this setting would benefit from more careful strategies, such as continued pretraining and balanced long-context data sources. Nonetheless, due to resource limitations, we directly pretrain the models on a 16K context length in this work, and hope that the results can reflect the ability of different architectures to handle longer contexts to some extent.
Due to the short web data used during pretraining, we consider evaluating the models on short-context tasks, or applying SFT on long-context tasks that contain real long sequences. As shown in Table 11, overall, the performance of RAT(L=16) stays in between of RNN and attention, and RAT(L=16)-SWA performs the best. It achieves a 0.02 pretraining lower perplexity than attention, and improves by about 1 point on average in commonsense reasoning and 0.8 points on QMSum. However, it underperforms attention around 0.7 points in F1 score on the NarrativeQA3 task, unlike in the 4K context length setting. On the other hand, the maximum throughput improves significantly: from 4.5$\sim$ 4.5$\times$ 4.5 × compared to attention at the 4K context window to 13.9$\sim$ 13.9$\times$ 13.9 × at 16K. This suggests that, to pursue higher accuracy performance, we may consider increasing the local attention window size (e.g., from 1024 to 4096) under the 16K context length setting.
C.4 Needle-in-Haystack (retrieval ability)
@@
To evaluate retrieval capabilities—an area where attention-based architectures are known to excel, and where RNNs typically fall short—we conduct experiments on the Needle-in-Haystack tasks. In these tasks, a “needle” (e.g., a magic number or UUID) is embedded within irrelevant passages, noisy text, or mixed with multiple key-value pairs. Each key-value pair is presented as a short identifier (the key) followed by its associated value, and the model is prompted to retrieve the correct value given a specific key.
@@
We adopt the RULER benchmark [55] and evaluate a range of Needle-in-Haystack task configurations. Specifically, task single_1 involves retrieving a specific number from a background filled with repeated noise. In single_2, the background is replaced with natural stories. single_3 increases the difficulty by requiring retrieval of a long and complex UUID. The multikey tasks are more challenging than the single-key ones, as they introduce multiple such key-value pairs into the context. In particular, multikey_2 and multikey_3 consist almost entirely of densely packed key-value pairs, among which only a single key is queried at the end. This makes the task especially challenging, as the model must retrieve the correct item from a highly cluttered input full of distractors. And multikey_3 further incorporates the complex UUID format. In the multiquery and multivalue settings, the model is required to resolve multiple retrieval targets—either by answering several distinct queries, or by retrieving all values associated with a single key. For detailed definitions of each task, we refer the reader to the benchmark documentation.
@@
During evaluation, we observed that models may fail to interpret certain prompts correctly. For instance, prompts like “A special magic number is hidden within the following text. Make sure to memorize it. I will quiz you about the number afterwards.” can cause failures even in attention-based models, especially on the harder multikey tasks. To mitigate this, we apply a light, one-round supervised fine-tuning stage before evaluation to adapt the models to the instruction patterns. We generate 1,000 synthetic training samples for each of the 8 tasks, resulting in a total of 8,000 examples disjoint from the validation sets. The models are trained on this dataset for one round and then are evaluated directly on all 8 tasks. This procedure yields a fairer and more stable comparison. As shown in Table 12, RNNs perform reasonably well on the simpler tasks (single_1, single_2), but their performance drops to near zero on the harder ones. Attention-based models perform consistently well, benefiting from their ability to access all previous tokens directly. RAT(L=16) achieves results close to full attention, especially in numeric tasks. However, it struggles on UUID tasks due to their complexity and the use of exact match scoring (single_3, multikey_3). Meanwhile, RAT(L=64) falls between RNNs and RAT(L=16), as expected given its partial access to long-range context.
@@
Table 12: Accuracy on the Needle-in-Haystack tasks with different configurations from the RULER benchmark [55], used to evaluate retrieval capabilities. Evaluation is conducted with a 4K context window, using exact match as the metric.
Model single_1 single_2 single_3 multikey_1 multikey_2 multikey_3 multiquery multivalue
RNN 99.4 99.8 0.0 2.2 0.0 0.0 5.8 5.7
RAT(L=64) 100.0 99.8 55.6 96.4 62.2 1.0 86.3 89.1
RAT(L=16) 99.6 100.0 94.6 99.6 99.6 82.6 91.2 94.8
Attention 100.0 100.0 100.0 99.6 99.6 95.2 98.6 99.1

### v3

@@ B.1 Latency
To supplement Subsec. 4.2 in main text, we put the concrete latency number of a single layer in Table 8, Table 9, and Table 10.
Table 8: Single temporal-mixing layer (including input and output projections) training time across different sequence lengths. The latency (ms) is tested on 200K tokens.
[…]
Table 9: Single temporal-mixing layer (including input and output projections) prefilling time across different sequences lengths. The latency (ms) is tested on 200K tokens.
[…]
Figure 4: We measure the maximum throughput of the full 1.3B model for generating 1024 tokens under different prefilling lengths. For each total sequence length $T$, the prefilling length is set to T1024T-1024. For example, $T=4096$ corresponds to a prefilling of 3072 tokens, while $T=8192$ and $T=16384$ correspond to 7168 and 15360 tokens, respectively. When the sequence length increases, the maximum throughput ratio between RAT and attention rises from $10.2\times$ to $15.6\times$, highlighting the strong efficiency advantage of RAT in long-context generation.
Table 10: Single temporal-mixing layer (including input and output projections) generation time at the specified position. The latency (ms) is tested on generating batches of tokens with $B=64$, $B=512$, and $B=1024$.
@@ B.2 Maximum throughput
As RAT also reduces cache memory usage, we report the maximum throughput of the full 1.3B model across different sequence lengths in Fig. 4. For example, RAT(L=16) achieves an approximately $10.2\times$ higher throughput than the baseline attention model at $T=4096$, and improves further to $15.6\times$ at $T=16384$. Similarly, RAT(L=16)-SWA reaches $4.5\times$, $8.0\times$, and $13.9\times$ higher throughput over attention at $T=4096$, 81928192, and $16384$, respectively.
@@ Appendix C Accuracy
C.1 Supplementary ablation study
All experiments here are conducted on the PG19 Book dataset with 200M-parameter models, and FLOPs are controlled for comparison.
Reversed design of RAT
Instead of employing recurrence within each chunk, we also explored a reversed design that applies attention inside the chunk, followed by recurrence to capture long-range dependencies. For this reversed design, standard RoPE is sufficient. Regarding the parameter allocation, we found it crucial to retain the key vector in the attention module as a full tensor, as sharing it together with the query vector leads to a collapse into single-head attention. Thus, the final reversed design shares the query vector across attention heads, employs low-rank matrices for the projection used for output gate, and preserves full projections for both the forget gate and the key vector.
Results of the reversed design are put in Table 11. It can be observed that, under identical FLOP constraints, this design significantly underperforms RAT. We think this is because short-context dependencies are easier to capture and do not suffer from memory degradation, so using lightweight recurrence locally is more efficient. Long‑context dependencies are harder and may require direct retrieval of distant information, where attention is more suitable. For example, to match the FLOPs budget of $\mathcal{O}(T/64)$, the reversed design requires chunk sizes of 128, 256, and 512 for sequence lengths of 8192, 16384, and 32768, respectively. With this small local window size, the long-range dependencies are attributed to the recurrence, thus leading to high perplexity. Only when we reduce the number of chunks to 16 does the perplexity drop below 14.00. Therefore, while the reversed design also allows for interpolation between attention and RNN (reducing to attention when $L=T$, and to RNN when $L=1$), we choose to focus on RAT due to its more efficient utilization of FLOPs.
Comparison to FLASH [gau]
As pointed out in Sec. 5, gau also uses the concept of chunking; however, we adopt a fundamentally different framework, both in the components involved and in how the two outputs are combined. Specifically, gau simply adds the outputs of inter-chunk and intra-chunk computations, whereas RAT organizes them hierarchically, applying recurrence over the key and value before inter-chunk attention. We argue that these differences lead to the weaker performance of gau, as shown in Table 13. First, as discussed above, RAT utilizes FLOPs more effectively than methods based on local attention, since its long-range dependencies must rely on memory-degrading components such as recurrence and linear attention. Second, we find that directly adding the outputs of softmax-based attention and linear attention introduces differences in output scale and potential representational conflicts, which may lower down the performance.
Comparison to dilated attention
We discuss the difference between RAT and the design that interleaves recurrence and dilated attention [longformer] with the sliding pattern. First, from an efficiency perspective, RAT has advantages over dilated attention in both training and inference. During training, dilated attention reduces parallelism by the dilation rate, since tokens within the same dilation span attend to different KV vectors. In contrast, RAT allows tokens to share the same chunk-level representation, except within their own chunk. During inference, the sliding dilation pattern causes a memory issue, since the KV cache of all tokens must be stored, whereas RAT reduces the cache size proportionally to the chunk size. Second, in terms of accuracy, to match the FLOPs, we set the dilation rate to 64 for RNN-Dilated attention, since dilated attention occurs in only half of the layers. It can be seen from Table 14 that dilated attention performs poorly. We observed significantly slower convergence at the beginning of training, likely due to its lack of global perception.
@@ C.1 Supplementary ablation study
Table 13: Performance of different chunk-based designs with $T=16384$ and FLOPs $\mathcal{O}(T/64)$.
Method Intra-chunk Inter-chunk Organization PPL RAT(L=64) Recurrence Attention Hierarchically 13.34 RAT-Reversed(C=64) Attention Recurrence Hierarchically 14.06 FLASH (C=128) Attention Linear attention Add together 14.4
Table 14: Performance compared to dilated attention with sliding patterns under the same FLOPs.
Method PPL RNN-Dilated attention 15.22 RAT(L=128) 13.44
@@ C.2 Length generalization
Because of the use of softmax-based attention and RoPE at the inter-chunk level, it is reasonable to expect that RAT may also face challenges in length generalization, where the model is trained on short sequences but evaluated on much longer ones. To study this, we consider SWA variants using either RoPE or NoPE for attention and RAT, following recent practices that interleave attention with local attention modules and apply positional encodings only in the local attention. This design has been adopted in practice and has shown promising results [cohere_llm].
@@
In conclusion, we find that RAT can outperform the attention module in length generalization, both with RoPE and with NoPE. Still, the problem is not fully solved, as none of the models reach the stability level of the simple RNN. While many techniques have been proposed to improve length extrapolation in attention models, they should also be considered for RAT, as RAT essentially redirects attention from every position to inter-chunk locations. For example, RoPE extrapolation in the attention module can be improved using methods from yarn, rope_ntk, and NoPE extrapolation has been studied in nope_head_scale, which points out that NoPE still has a context length limit, though it performs better than RoPE. Their work links the failure to shifts in attention distributions and proposes tuning the temperature of attention heads to improve extrapolation. We believe such methods could also be applied to RAT, and leave further exploration of these techniques to future work.
C.3 Needle-in-Haystack (retrieval ability)
@@
To evaluate retrieval capabilities—an area where attention-based architectures are known to excel, and where RNNs typically fall short—we conduct experiments on the Needle-in-Haystack tasks. In these tasks, a "needle" (e.g., a magic number or UUID) is embedded within irrelevant passages, noisy text, or mixed with multiple key-value pairs. Each key-value pair is presented as a short identifier (the key) followed by its associated value, and the model is prompted to retrieve the correct value given a specific key.
@@
We adopt the RULER benchmark [ruler] and evaluate a range of Needle-in-Haystack task configurations. Specifically, task single_1 involves retrieving a specific number from a background filled with repeated noise. In single_2, the background is replaced with natural stories. single_3 increases the difficulty by requiring retrieval of a long and complex UUID. The multikey tasks are more challenging than the single-key ones, as they introduce multiple such key-value pairs into the context. In particular, multikey_2 and multikey_3 consist almost entirely of densely packed key-value pairs, among which only a single key is queried at the end. This makes the task especially challenging, as the model must retrieve the correct item from a highly cluttered input full of distractors. And multikey_3 further incorporates the complex UUID format. In the multiquery and multivalue settings, the model is required to resolve multiple retrieval targets, either by answering several distinct queries or by retrieving all values associated with a single key. For detailed definitions of each task, we refer the reader to the benchmark documentation.
@@
During evaluation, we observed that models may fail to interpret certain prompts correctly. For instance, prompts like "A special magic number is hidden within the following text. Make sure to memorize it. I will quiz you about the number afterwards." can cause failures even in attention-based models, especially on the harder multikey tasks. To mitigate this, we apply a light, one-round supervised fine-tuning stage before evaluation to adapt the models to the instruction patterns. We generate 1000 synthetic training samples for each of the 8 tasks, resulting in a total of 8000 examples disjoint from the validation sets. The models are trained on this dataset for one round and then are evaluated directly on all 8 tasks. This procedure yields a fairer and more stable comparison. As shown in Table 6, RNNs perform reasonably well on the simpler tasks (single_1, single_2), but their performance drops to near zero on the harder ones. Attention-based models perform consistently well, benefiting from their ability to access all previous tokens directly. RAT(L=16) achieves results close to full attention, especially in numeric tasks. However, it struggles on UUID tasks due to their complexity and the use of exact match scoring (single_3, multikey_3). Meanwhile, RAT(L=64) falls between RNNs and RAT(L=16), as expected given its partial access to long-range context.